# Effective (Floquet) Lindblad generators from spectral unwinding

### v1

@@ Abstract
A mathematical description of the reduced dynamics of an open quantum system can often be given in terms of a completely positive and trace preserving (CPTP) map, also known as quantum channel. In a seminal work by Wolf et al. [Phys. Rev. Lett. 101, 150402 (2008)], it was shown that deciding whether a given quantum channel was generated from an underlying effective Markovian dynamics, with time-independent Lindbladian generator, is generally an NP-hard problem. In this work we show that in cases where one has access to the full reduced dynamics at all previous times (the dynamical map) one can significantly facilitate the search for an effective generator by making use of Floquet theory. By performing a spectral unwinding such that the effective micromotion is minimized, the effective Floquet generator is often an excellent candidate for an effective Lindblad generator, hence significantly reducing the complexity of the search for an effective Lindblad generator in many (though not all) cases. Our results are relevant for engineering Floquet Lindbladians in complex many-body systems.
@@
Introduction.—Digital and analog simulation of quantum many-body systems have yielded fascinating results [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18]. Both rely in part on the advent of Floquet engineering [19, 20, 5, 6, 21] (which in the context of digital simulation is usually referred to as trotterization [22, 23, 24, 25, 26, 27, 28, 29]). The central idea is that for an isolated quantum system with time-periodic Hamiltonian $H(t)=H(t+T)$, the one-cycle time-evolution operator $U(T)$ can be rewritten as $U(T)=\exp(-iH_{\mathrm{F}}T)$ with a time-independent Floquet Hamiltonian $H_{\mathrm{F}}$ with possibly novel terms that are not present in the lab Hamiltonian, e.g. artificial gauge fields [3, 4, 30, 19, 5, 6, 31, 32, 33] or nontrivial interactions [34, 35, 36, 37, 13, 38, 39].
@@
Recently, the question was raised [40, 41, 42] whether these ideas can be generalized to open Floquet systems [43], i.e. that interact with an environment (e.g. by coupling the system to external photon or particle reservoirs). Interestingly, it was shown [41] that the dynamics under a time-periodic Lindbladian $\mathcal{L}(t)=\mathcal{L}(t+T)$ [41, 27, 44, 45, 46, 47] cannot always be understood as generated from an effective time-independent Floquet Lindbladian $\mathcal{L}_{\mathrm{F}}$ [40, 48, 49, 41, 42, 50, 51, 44, 52]. This is due to effective non-Markovian effects that are built up during the evolution. Nevertheless, as a result of the existence of branches when calculating the complex logarithm of the dynamical map, deciding whether a valid Floquet Lindbladian exists is generally an NP-hard problem [53]. This question closely relates to earlier work by Wolf et al. [54, 53] where the existence of an effective Lindbladian generator was studied for a general completely positive and trace-preserving (CPTP) map $\mathcal{V}$. In this work we show that, by using ideas from Floquet theory, the general problem of deciding Markovianity can be addressed more efficiently if the full dynamical map $\mathcal{V}(t)$ is known for all times $t$. Our results lay the foundation for Floquet engineering of Lindbladians in extended many-body systems.
@@
Quantum Channels, Dynamical maps.—The mathematical cornerstone of the theory of Markovian open quantum systems are one-parameter semigroups $\mathcal{V}(t)\in L(L(\mathcal{H}))$, where $L(\mathcal{H})$ denotes the space of linear operators on the Hilbert space $\mathcal{H}$ with dimension $d$. The semigroup $\mathcal{V}(t)=e^{\mathcal{L}t}$, $t\geq 0$, describes an evolution of the density operator $\varrho(t)=\mathcal{V}(t)[\varrho(0)]$ that is Markovian and consistent with a physical evolution, if the dynamical map $\mathcal{V}(t)$ is CPTP, i.e. a quantum channel, at all times $t$. Lindblad [55], Gorini, Kossakowski and Sudarshan [56] have shown that a superoperator $\mathcal{L}$ is the generator of a quantum dynamical semigroup $\mathcal{V}(t)=e^{\mathcal{L}t}$, iff it can be expressed as
@@
Markovianity test.—Let us turn to the converse question and ask “Given an arbitrary quantum channel $\mathcal{V}$ (say at time $T$) is there an effective generator $\mathcal{S}=\log(\mathcal{V})/T$ that has Lindblad form as in Eq. (1)?”. Due to the multi-valuedness of the complex logarithm, this question turns out to result in an NP-hard problem in general, also called the Markovianity problem [54, 53]. An algorithm to address this problem was proposed by Wolf et al. [54, 53]: A generator $\mathcal{S}\in L(L(\mathcal{H}))$ has Lindblad form, iff (i) for all Hermitian $\kappa=\kappa^{\dagger}$, $\mathcal{S}$ preserves Hermiticity, i.e. $\mathcal{S}[\kappa]=\bigl{(}\mathcal{S}[\kappa]\bigr{)}^{\dagger}$, and (ii) $\mathcal{S}$ is conditionally completely positive (CCP), i.e.
@@
Here, $\Sigma={|{\Omega}\rangle}{\langle{\Omega}|}$ is the projector onto the maximally entangled state ${|{\Omega}\rangle}=\frac{1}{\sqrt{d}}\sum_{j=1}^{d}{|{j}\rangle}\otimes{|{j}\rangle}$ of the system and an ancilla of same dimension $d$. $\mathcal{S}^{\Gamma}=d\cdot(\mathcal{S}\hskip 1.42262pt\otimes\hskip 1.42262pt\mathbf{1}_{d})[\Sigma]$ denotes the Choi representation of $\mathcal{S}$. Since the complex logarithm is not uniquely defined, we need to check if there is a branch of the operator logarithm of $\mathcal{V}$ that satisfies condition (i) and (ii).
@@
To do this, we introduce the scalar product $\langle\!\langle A|B\rangle\!\rangle=\operatorname{Tr}[A^{\dagger}B]$ on the Hilbert space $L(\mathcal{H})$. By also defining an orthonormal basis $\{F_{\alpha}\}_{\alpha=1}^{d^{2}}$ in $L(\mathcal{H})$, we can represent quantum channels $\mathcal{V}$ as matrices $\hat{\mathcal{V}}$, with their corresponding matrix elements $\hat{\mathcal{V}}_{\alpha\beta}={\langle\!\langle{F_{\alpha}}|}\mathcal{V}{|{F_{\beta}}\rangle\!\rangle}$. On this space, (density) matrices $\varrho={|{i}\rangle}{\langle{j}|}$ become vectors ${|{\varrho}\rangle\!\rangle}={|{i,j}\rangle\!\rangle}$, hence this identification is known as vectorization of the density matrix. We first cast the map $\hat{\mathcal{V}}$ into Jordan normal form [54], $\hat{\mathcal{V}}=\sum_{r}\lambda_{r}{|{r_{r}}\rangle\!\rangle}{\langle\!\langle{l_{r}}|}+\sum_{c=1}^{N_{c}}\bigl{(}\lambda_{c}{|{r_{c}}\rangle\!\rangle}{\langle\!\langle{l_{c}}|}+\lambda_{c}^{*}\ \mathbb{F}{|{r_{c}}\rangle\!\rangle}{\langle\!\langle{l_{c}}|}\mathbb{F}\bigr{)}$, with $r$ indexing the real and $c$ the complex eigenvalues and ${N_{c}}$ the number of complex conjugated pairs. The operator $\mathbb{F}$ denotes the flip-operator, $\mathbb{F}(\sum_{i,j}c_{ij}{|{i,j}\rangle\!\rangle})=\sum_{i,j}c_{ij}^{*}{|{j,i}\rangle\!\rangle}$, which relates the spectral projectors of complex conjugated eigenvalue pairs. By taking the logarithm, we obtain a family of possible generators of $\mathcal{V}$ [54, 41],
$\hat{\mathcal{S}}_{\{\vec{x}\}}=\hat{\mathcal{S}}_{0}+i\frac{2\pi}{T}\sum_{c=1}^{N_{c}}x_{c}\left({|{r_{c}}\rangle\!\rangle}{\langle\!\langle{l_{c}}|}-\mathbb{F}{|{r_{c}}\rangle\!\rangle}{\langle\!\langle{l_{c}}|}\mathbb{F}\right),$ (3)
@@
with $\hat{\mathcal{S}}_{0}$ stemming from the principal branch of the logarithm and a vector $\{\vec{x}\}=\{x_{1},x_{2},...,x_{N_{c}}\}\in\mathbb{Z}^{N_{c}}$ of integers labeling the possible branches of the complex logarithm. If all real eigenvalues $\lambda_{r}$ of $\hat{\mathcal{V}}$ are positive, all of these generators $\mathcal{S}_{\{\vec{x}\}}$ preserve hermiticity by construction and hence fulfil condition (i). At first glance, to test condition (ii), we have to inspect all branches, i.e., a countably infinite number of combinations of $N_{c}$ integers. Nevertheless, by use of methods from integer programming the problem can be reduced to a smaller set of possible integers [53], or alternatively one can apply machine-learning algorithms [57].
@@
However it was shown [53] that deciding Markovianity is generally an NP-hard problem. This is especially detrimental in the case of interacting many-body systems where already the underlying Hilbert space $\mathcal{H}$ grows exponentially with system size $L$. Consider for example a spin chain of length $L$. Then, $\mathrm{dim}(\mathcal{H})=2^{L}$, so the maximum number of complex pairs (since $\lambda_{r}=1$ always exists [58]) is given by $N_{c}\leq 2^{2L-1}-1$. If we would just check the two closest branches around the principle branch, i.e. $x_{c}\in\{-1,0,1\}$, we already find super exponential scaling $3^{2^{2L-1}-1}$ of the number of branches for which condition (ii) has to be checked.
@@
In the case that no branch gives rise to a valid Lindbladian, one can measure the distance to Markovianity and find the branch that is closest to a Markovian evolution. A possible measure proposed by Wolf et al. [54] that we use is based on adding a noise term $\chi\mathcal{Z}$ of strength $\chi$ to the generator $\mathcal{S}_{\{\vec{x}\}}$. The operator $\mathcal{Z}$ is the generator of the depolarizing channel $\exp({T\chi\mathcal{Z}})[\varrho]=e^{-\chi T}\varrho+(1-e^{-\chi T})\mathbf{1}/d$. The distance from Markovianity is defined as [54]
@@
the minimal strength needed, such that the generator $\mathcal{S}_{\{\vec{x}\}}+\chi\mathcal{Z}$ is Lindbladian.
Systems with access to full time-dependant map.—In the following we describe and test a simple strategy that allows to address (and often solve) this problem in cases, where additionally to $\mathcal{V}$ at time $T$, the full dynamical map $\mathcal{V}(t)$ is known also for times $t\in[0,T]$. To this end, we make connection to Floquet theory by continuing $\mathcal{V}(t)$ periodically for $t>T$, e.g. $\mathcal{V}(t)=\mathcal{V}(t-T)\mathcal{V}(T)$ for $T<t\leq 2T$ and so on. Then, the time evolution can effectively understood as generated from a time-periodic generator $\mathcal{G}(t)=\mathcal{G}(t+T)$ (not necessarily of Lindblad form) since
@@
Let us briefly discuss Floquet theory for isolated systems. For systems with time-periodic Hamiltonian $H(t)=H(t+T)$, the fundamental solutions can be written in terms of Floquet states ${|{\psi_{\alpha}(t)}\rangle}=\exp(-i\varepsilon_{\alpha}t){|{u_{\alpha}(t)}\rangle}$ with time-periodic Floquet modes ${|{u_{\alpha}(t)}\rangle}={|{u_{\alpha}(t+T)}\rangle}$ and quasienergies $\varepsilon_{\alpha}$ [20]. Stroboscopically, at $t=0,T,2T,\dots$ the time evolution operator $U(T)=\exp(-iH_{\mathrm{F}}T)$ is described by the effective time-independent Floquet Hamiltonian $H_{\mathrm{F}}=\sum_{\alpha}\varepsilon_{\alpha}{|{u_{\alpha}(0)}\rangle}{\langle{u_{\alpha}(0)}|}$. Note that the quasienergies $\varepsilon_{\alpha}$ are not uniquely defined and can be redefined under the gauge transformation $\varepsilon_{\alpha}\rightarrow\varepsilon_{\alpha}+m\omega$ and ${|{u_{\alpha}(t)}\rangle}\rightarrow e^{im\omega t}{|{u_{\alpha}(t)}\rangle}$, where $\omega=2\pi/T$. Note that picking a specific gauge is equivalent to fixing the branch of the complex logarithm when solving for $H_{\mathrm{F}}=i\log[U(T)]/T$.
We now apply Floquet’s theorem to Eq. (5). The Floquet states take the form ${|{\varrho_{\mu}(t)}\rangle\!\rangle}=e^{-i\Omega_{\mu}t}{|{\Phi_{\mu}(t)}\rangle\!\rangle}$, where $\Omega_{\mu}$ are complex in general and the Floquet modes ${|{\Phi_{\mu}(t)}\rangle\!\rangle}$ are periodic in time [42, 58, 20]. Note that also here, redefining $\Omega_{\mu}\rightarrow\Omega_{\mu}+m\omega$ and ${|{\Phi_{\mu}(t)}\rangle\!\rangle}\rightarrow e^{im\omega t}{|{\Phi_{\mu}(t)}\rangle\!\rangle}$, with $m\in\mathbb{Z}$, leaves ${|{\varrho_{\mu}(t)}\rangle\!\rangle}$ invariant. In case of a purely coherent evolution, $\mathcal{G}(t)=-i[H(t),\cdot]$, the $\Omega_{\mu}$ correspond to quasienergy differences. In this paper, even in the case of dissipative dynamics, we still refer to the $\Omega_{\mu}$ as quasienergy differences, which are complex in general. For a given time $t$, we can decompose the dynamical map $\hat{\mathcal{V}}(t)$ as $\hat{\mathcal{V}}(t)=\sum_{\mu}e^{-i\Omega_{\mu}t}{|{\Phi_{\mu}(t)}\rangle\!\rangle}{\langle\!\langle{\tilde{\Phi}_{\mu}(0)}|}$, with the left Floquet modes ${\langle\!\langle{\tilde{\Phi}_{\mu}(t)}|}$ being different from the right ones for dissipative systems with a non-hermitian generator [42]. In this way, we can immediately identify the effective generator as [42] $\hat{\mathcal{S}}=-\sum_{\mu}i\Omega_{\mu}{|{\Phi_{\mu}(T)}\rangle\!\rangle}{\langle\!\langle{\tilde{\Phi}_{\mu}(0)}|},$ with the freedom of choosing the quasienergy differences $\Omega_{\mu}$ representing the freedom of choosing a branch of the logarithm in Eq. (3). Henceforth, we write ${|{\Phi_{\mu}(T)}\rangle\!\rangle}={|{\Phi_{\mu}(0)}\rangle\!\rangle}\equiv{|{\Phi_{\mu}}\rangle\!\rangle}$ and ${\langle\!\langle{\tilde{\Phi}_{\mu}(0)}|}\equiv{\langle\!\langle{\tilde{\Phi}_{\mu}}|}$.
@@
Figure 1: (a)-(f) Distance from Markovianity $\mu$ (which here equals to distance from von-Neumann form $\mathcal{N}$) of the effective generator of the one-cycle evolution superoperator as a function of driving strength $E$ and frequency $\omega$, for no dissipation, $\gamma=0$, coupling strength $w=0.1$, and chain lengths of (a),(b),(d),(e) $L=2$ and (c),(f) $L=3$. To lift degeneracies in the spectrum, we set $\Delta_{1}=1.01,\Delta_{2}=1.00,\Delta_{3}=0.98$.
@@
Let us now, for illustrative purposes regard the evolution with a time-independent generator $\partial_{t}\varrho(t)=\mathcal{L}\varrho(t)$, such that $\mathcal{V}(T)=e^{\mathcal{L}T}$, as an effective time-periodic problem with arbitrary period $T$. If we choose a large $T$ such that $\omega=2\pi/T$ is smaller than the imaginary part of some of the eigenvalues of $\mathcal{L}$, we observe that the generator $\mathcal{S}_{0}$ that we obtain from the principle branch of $\log(\mathcal{V}(T))/T$ is not identical to $\mathcal{L}$ but there occurs a “winding” of the eigenvalues of the Floquet generator $\mathcal{S}$ into the first Floquet Brillouin zone, $\mathrm{Re}\Omega_{\mu}\in[-\omega/2,\omega/2)$. This winding turns out to make the previously outlined procedure of searching around the principle branch inefficient, even for a fully coherent time evolution with no dissipation: To demonstrate this, we consider a circularly driven spin chain of length $L$

### v3

@@ Abstract
A mathematical description of the reduced dynamics of an open quantum system can often be given in terms of a completely positive and trace preserving (CPTP) map, also known as quantum channel. In a seminal work by Wolf et al. [Phys. Rev. Lett. 101, 150402 (2008)], it was shown that deciding whether a given quantum channel was generated from an underlying effective Markovian dynamics, with time-independent generator of Lindblad form, is generally an NP-hard problem. The difficulty is related to the fact that one has to search through all possible branches of the operator logarithm of the map, in order to identify, if any of the resulting effective generators is of Lindblad form. In this work we show that in cases where one has access to the full reduced dynamics at all previous times (the dynamical map) one can significantly facilitate the search for an effective generator by making use of Floquet theory. By performing a spectral unwinding such that the effective micromotion is minimized, the effective Floquet generator is often an excellent candidate for an effective generator of Lindblad form. This significantly reduces the complexity of the search for an effective Lindbladian in many (though not all) cases. Our results are relevant for engineering Floquet Lindbladians in complex many-body systems.
I Introduction
@@
Digital and analog simulation of quantum many-body systems in engineered quantum systems based on photons or atoms have yielded fascinating results [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18]. Both rely in part on the advent of Floquet engineering [19, 20, 5, 6, 21] (which in the context of digital simulation is usually referred to as trotterization [22, 23, 24, 25, 26, 27, 28, 29]). The central idea is that for an isolated quantum system with time-periodic Hamiltonian $H(t)=H(t+T)$, the one-cycle time-evolution operator $U(T)$ can be rewritten as $U(T)=\exp(-iH_{\mathrm{F}}T)$ with a time-independent Floquet Hamiltonian $H_{\mathrm{F}}$ with possibly novel terms that are not present in the lab Hamiltonian, e.g. artificial gauge fields [3, 4, 30, 19, 5, 6, 31, 32, 33] or nontrivial interactions [34, 35, 36, 37, 13, 38, 39].
@@
Recently, the question was raised [40, 41, 42] whether these ideas can be generalized to open Floquet systems [43], i.e. systems that interact with an environment (e.g. by coupling the system to external photon or particle reservoirs). Interestingly, it was shown [41] that the dynamics under a time-periodic Lindbladian $\mathcal{L}(t)=\mathcal{L}(t+T)$ [41, 27, 44, 45, 46, 47] cannot always be understood as generated from an effective time-independent Floquet Lindbladian $\mathcal{L}_{\mathrm{F}}$ [40, 48, 49, 41, 42, 50, 51, 44, 52]. This is due to effective non-Markovian effects that are built up during the evolution. As a result of the existence of branches when calculating the complex logarithm of the dynamical map, deciding whether a valid Floquet Lindbladian exists is generally an NP-hard problem [53]. This question closely relates to earlier work by Wolf et al. [54, 53] where the existence of an effective Lindbladian generator was studied for a snapshot in time of a general quantum evolution $\mathcal{V}$, a completely positive and trace-preserving (CPTP) map, irrespective of where it originates from. In this work, we show that by using ideas from Floquet theory, the general problem of deciding Markovianity can be addressed more efficiently in cases where one does not only have access to such a snapshot in time, but also to the full dynamical map $\mathcal{V}(t)$ for all times $t$. This is relevant, e.g. for Floquet engineering of open quantum systems, where such effective Lindbladian generators can capture the system’s stroboscopic dynamics and hence be utilized in the same manner as Floquet Hamiltonians in the case of isolated systems. Our results lay the foundation for Floquet engineering of Lindbladians in extended many-body systems.
II Markovianity test
@@
The mathematical cornerstone of the theory of Markovian open quantum systems are one-parameter semigroups $\mathcal{V}(t)$ which are superoperators on the Hilbert space $\mathcal{H}$ with dimension $d$. The semigroup $\mathcal{V}(t)=e^{\mathcal{L}t}$, $t\geq 0$, describes an evolution of the density operator $\varrho(t)=\mathcal{V}(t)[\varrho(0)]$ that is Markovian and consistent with a physical evolution, if the dynamical map $\mathcal{V}(t)$ is CPTP, i.e. a quantum channel, at all times $t$. Lindblad [55], Gorini, Kossakowski and Sudarshan [56] have shown that a superoperator $\mathcal{L}$ is the generator of a quantum dynamical semigroup $\mathcal{V}(t)=e^{\mathcal{L}t}$, iff it can be expressed as
@@
Let us turn to the converse question and ask “Given an arbitrary quantum channel $\mathcal{V}$ (say at time $T$) is there an effective generator
$\mathcal{S}=\log(\mathcal{V})/T$ (2)
that has Lindblad form as in Eq. (1)?”. Due to the multi-valuedness of the complex logarithm, this question turns out to result in an NP-hard problem in general, also called the Markovianity problem [54, 53]. A criterion for Markovianity in this sense was proposed by Wolf et al. [54, 53]: A generator $\mathcal{S}$ has Lindblad form, iff (i) for all Hermitian $\kappa=\kappa^{\dagger}$, $\mathcal{S}$ preserves Hermiticity, i.e. $\mathcal{S}[\kappa]=\bigl{(}\mathcal{S}[\kappa]\bigr{)}^{\dagger}$, and (ii) $\mathcal{S}$ is conditionally completely positive (CCP), i.e.
@@
Here, $\Sigma={|{\Omega}\rangle}{\langle{\Omega}|}$ is the projector onto the maximally entangled state ${|{\Omega}\rangle}=\frac{1}{\sqrt{d}}\sum_{j=1}^{d}{|{j}\rangle}\otimes{|{j}\rangle}$ of the system and an ancilla of same dimension $d$ and
$\mathcal{S}^{\Gamma}=d\cdot(\mathcal{S}\hskip 1.42262pt\otimes\hskip 1.42262pt\mathbf{1}_{d})[\Sigma]$ (4)
denotes the Choi representation of $\mathcal{S}$. Since the complex logarithm is not uniquely defined, we need to check if there is a branch of the operator logarithm of $\mathcal{V}$ that satisfies condition (i) and (ii).
@@
To do this, we introduce the operator scalar product $(A|B)=\operatorname{Tr}[A^{\dagger}B]$. By also defining an orthonormal operator basis $\{F_{\alpha}\}_{\alpha=1}^{d^{2}}$, we can represent quantum channels $\mathcal{V}$ as matrices $\hat{\mathcal{V}}$, with their corresponding matrix elements $\hat{\mathcal{V}}_{\alpha\beta}={({F_{\alpha}}|}\mathcal{V}{|{F_{\beta}})}$, which allows for the identification $\varrho={|{i}\rangle}{\langle{j}|}\rightarrow{|{\varrho})}={|{i,j})}$, known as vectorization of (density) matrices. A useful operation is $\mathbb{F}\bigl{[}\sum_{i,j}c_{ij}{|{i,j})}\bigr{]}=\sum_{i,j}c_{ij}^{*}{|{j,i})}$ [54]. Assuming the map $\hat{\mathcal{V}}$ to be non-defective, which is generally justified for Markovian maps and quantum channels [53], we first cast it into Jordan normal form [54],
$\hat{\mathcal{V}}=\sum_{r}\lambda_{r}{|{r_{r}})}{({l_{r}}|}+\sum_{c=1}^{N_{c}}\bigl{[}\lambda_{c}{|{r_{c}})}{({l_{c}}|}+\lambda_{c}^{*}\ \mathbb{F}{|{r_{c}})}{({l_{c}}|}\mathbb{F}\bigr{]},$ (5)
with $r$ and $c$ indexing the real and complex eigenvalues, respectively, and ${N_{c}}$ the total number of complex conjugated pairs. The logarithm yields a family of possible generators of $\mathcal{V}$ [54, 41],
$\hat{\mathcal{S}}_{\vec{x}}=\hat{\mathcal{S}}_{0}+i\frac{2\pi}{T}\sum_{c=1}^{N_{c}}x_{c}\Bigl{[}{|{r_{c}})}{({l_{c}}|}-\mathbb{F}{|{r_{c}})}{({l_{c}}|}\mathbb{F}\Bigr{]},$ (6)
@@
with $\hat{\mathcal{S}}_{0}$ stemming from the principal branch of the logarithm and a vector $\vec{x}=\{x_{1},x_{2},...,x_{N_{c}}\}\in\mathbb{Z}^{N_{c}}$ of integers labeling the possible branches of the complex logarithm. Note that, to arrive at Eq. (6), we implicitly assume that condition (i) is fulfilled [54, 41], i.e. that all real eigenvalues $\lambda_{r}$ of $\hat{\mathcal{V}}$ are positive  [^1] To fulfil condition (i), negative real eigenvalues $\lambda_{r}<0$ can only occur with an even degeneracy, such that one can reinterpret them as complex conjugated pairs. Naively, to test condition (ii), we have to inspect all branches, i.e., a countably infinite number of combinations of $N_{c}$ integers. Nevertheless, by use of methods from integer programming, the problem can be reduced to a smaller set of possible integers [53], or alternatively one can apply machine-learning algorithms [58].
@@
However, it was shown [53] that deciding Markovianity is generally an NP-hard problem. This is especially detrimental in the case of interacting many-body systems where already the underlying Hilbert space $\mathcal{H}$ grows exponentially with system size $L$. Consider for example a spin chain of length $L$, with $d=\mathrm{dim}(\mathcal{H})=2^{L}$. If we would just naively check the two closest branches around the principle branch, i.e. $x_{c}\in\{-1,0,1\}$, we already find super exponential scaling $3^{2^{2L-1}-1}$ of the number of branches for which condition (ii) has to be checked.
@@
In the case that no branch gives rise to a valid Lindbladian, one can measure the distance to Markovianity and find the branch that is closest to a Markovian evolution. We use a measure proposed by Wolf et al. [54] that is based on adding a noise term $\chi\mathcal{Z}$ of strength $\chi$ to the generator $\mathcal{S}_{\{\vec{x}\}}$, where $\mathcal{Z}$ is the generator of the depolarizing channel. The distance from Markovianity is defined as [54]
@@
i.e. the minimal strength needed, such that the generator $\mathcal{S}_{\{\vec{x}\}}+\chi\mathcal{Z}$ is Lindbladian. Details on the definition and how to calculate this measure $\mu$ are laid out in Refs. [54, 41].
III Systems with access to full time-dependent map
We describe a strategy that allows to address (and often solve) this problem in cases, where additionally to $\mathcal{V}$ at time $T$, the full dynamical map $\mathcal{V}(t)$ is known also for times $t\in[0,T]$. To this end, we draw a connection to Floquet theory by continuing $\mathcal{V}(t)$ periodically for $t>T$, e.g. $\mathcal{V}(t)=\mathcal{V}(t-T)\mathcal{V}(T)$ for $T<t\leq 2T$ and so on (a similar strategy was employed in the context of isolated systems for the purpose of adiabatically preparing eigenstates of time evolution operators [59]). Then, the time evolution can effectively be understood as generated from a time-periodic generator $\mathcal{G}(t)=\mathcal{G}(t+T)$ (not necessarily of Lindblad form) since
@@ III Systems with access to full time-dependent map
Let us briefly discuss Floquet theory for isolated (and open) systems. For systems with time-periodic Hamiltonian $H(t)=H(t+T)$ (generator $\mathcal{G}(t)$), the fundamental solutions can be written in terms of Floquet states ${|{\psi_{\alpha}(t)}\rangle}=\exp(-i\varepsilon_{\alpha}t){|{u_{\alpha}(t)}\rangle}$ $\bigl{(}{|{\varrho_{\mu}(t)})}=\exp(-i\Omega_{\mu}t){|{\Phi_{\mu}(t)})}\bigr{)}$ with time-periodic Floquet modes ${|{u_{\alpha}(t)}\rangle}={|{u_{\alpha}(t+T)}\rangle}\ \bigl{(}{|{\Phi_{\mu}(t)})}={|{\Phi_{\mu}(t+T)})}\bigr{)}$ [42, 60, 20] and quasienergies $\varepsilon_{\alpha}$ [20] (eigenvalues $\Omega_{\mu}$). In case of a purely coherent evolution, $\mathcal{G}(t)=-i[H(t),\cdot]$, the $\Omega_{\mu}$ correspond to quasienergy differences (a term that we will continue to use also for the dissipative case). Note that the quasienergies $\varepsilon_{\alpha}$ (quasienergy differences $\Omega_{\mu}$) are not uniquely defined and can be redefined under the gauge transformation $\varepsilon_{\alpha}\rightarrow\varepsilon_{\alpha}+m\omega$ and ${|{u_{\alpha}(t)}\rangle}\rightarrow e^{im\omega t}{|{u_{\alpha}(t)}\rangle}\ \bigl{(}\Omega_{\mu}\rightarrow\Omega_{\mu}+m\omega$ and ${|{\Phi_{\mu}(t)})}\rightarrow e^{im\omega t}{|{\Phi_{\mu}(t)})}\bigr{)}$, where $\omega=2\pi/T$ and $m\in\mathbb{Z}$. In the isolated case, picking a specific gauge is equivalent to fixing the branch of the complex logarithm when solving for $H_{\mathrm{F}}=i\log[U(T)]/T$. Stroboscopically, at $t=0,T,2T,\dots$ the time evolution operator $U(T)=\exp(-iH_{\mathrm{F}}T)$ is described by the effective time-independent Floquet Hamiltonian
$H_{\mathrm{F}}=\sum_{\alpha}\varepsilon_{\alpha}{|{u_{\alpha}(0)}\rangle}{\langle{u_{\alpha}(0)}|}.$ (9)
In the dissipative case, for a given time $t$, we can decompose the dynamical map $\hat{\mathcal{V}}(T)$ as
$\hat{\mathcal{V}}(T)=\sum_{\mu}e^{-i\Omega_{\mu}T}{|{\Phi_{\mu}(T)})}{({\tilde{\Phi}_{\mu}(0)}|},$ (10)
with the left Floquet modes ${({\tilde{\Phi}_{\mu}(t)}|}$ being different from the right ones for dissipative systems with a non-hermitian generator [42]. In this way, we can immediately identify the effective generator as [42]
$\hat{\mathcal{S}}=-\sum_{\mu}i\Omega_{\mu}{|{\Phi_{\mu}(0)})}{({\tilde{\Phi}_{\mu}(0)}|},$ (11)
with the freedom of choosing the Floquet gauge of the quasienergy differences $\Omega_{\mu}$ indicating the freedom of choosing a branch of the logarithm in Eq. (6).
@@
Figure 1: (a)-(f) Distance from Markovianity $\mu$ (which here equals to distance from von Neumann form $\mathcal{N}$) of the effective generator of the one-cycle evolution superoperator as a function of driving strength $E$ and frequency $\omega$, for no dissipation, $\gamma=0$, coupling strength $w=0.1$, and chain lengths of (a),(b),(d),(e) $L=2$ and (c),(f) $L=3$. To avoid numerical artifacts stemming from degeneracies in the spectrum of $\mathcal{V}(T)$, we set $\Delta_{1}=1.01,\Delta_{2}=1.00,\Delta_{3}=0.98$.
@@
Let us now, for illustrative purposes, regard the evolution with a time-independent generator, $\partial_{t}\varrho(t)=\mathcal{L}\varrho(t)$, such that $\mathcal{V}(T)=\exp({\mathcal{L}T})$, as an effective time-periodic problem with arbitrary period $T$. If we choose a large $T$ such that $\omega=2\pi/T$ is smaller than the imaginary part of some of the eigenvalues of $\mathcal{L}$, we observe that the generator $\mathcal{S}_{0}$ that we obtain from the principle branch of $\log(\mathcal{V}(T))/T$ is not identical to $\mathcal{L}$ but there occurs a “winding” of the eigenvalues of the Floquet generator $\mathcal{S}$ “around” the first Floquet Brillouin zone, $\mathrm{Re}[\Omega_{\mu}]\in[-\omega/2,\omega/2)$, since $\exp(-i\Omega_{\mu}T)$ has the shape of a circle. This winding turns out to make the previously outlined procedure of searching around the principle branch inefficient, even for a fully coherent time evolution with no dissipation: To demonstrate this, we consider a circularly driven spin chain of length $L$